# An Efficient Difference-of-Convex Solver for Privacy Funnel

### v1

@@ I Introduction
where $\eta_{X}$ denotes a control threshold. In the above problem, the joint distribution $P(X,Y)$ is assumed to be known and the random variables satisfy the Markov chain $Y\rightarrow X\rightarrow Z$. Here, $Y$ is the private information; $X$ the public information and $Z$ is the released (compressed) codes; $I(Z;X)$ denotes the mutual information between two random variables $Z$ and $X$. In (1), the variables to optimize with are the stochastic mappings $P(Z|X)$ with $\Omega_{Z|X}$ denotes the associated feasible solution set. The aim of the problem is to characterize the fundamental trade-off between the privacy leakage and the utility of the released information. One of the challenges in solving (1) is the non-convexity of the problem. It is well-known that the mutual information $I(Z;X)$ with known marginal $P(X)$ is a convex function with respect to $P(Z|X)$ [2]; Additionally, due to the Markov relation, $I(Z;Y)$ is a convex function of $P(Z|X)$ as well. Together, (1) minimizes a convex objective under a non-convex set. Previous solvers for PF address the non-convexity by restricting the feasible solution set further for efficient computation. Both [1] and [3] consider discrete settings and focus on deterministic $P(z|x)$, i.e., a realization of $x\in\mathcal{Z}$ belongs to a single cluster $z\in\mathcal{Z}$ only. They propose greedy algorithms to solve the reduced problem and differ in choosing the subset of $\mathcal{Z}$ to combine with. However, the imposed restriction limits the characterization of the privacy-utility trade-off. In our earlier work [4], we propose a non-convex alternating direction methods of multipliers (ADMM) solver for PF which is non-greedy, but it suffers from slow convergence [5].
@@
Another challenge of PF is knowing the joint distribution $P(X,Y)$. Overcoming this is important for practical application of the method as $P(X,Y)$ is prohibitively difficult to obtain in general. Nonetheless, it is relatively easy to obtain samples of the joint distribution, which become a training dataset as an indirect access to $P(X,Y)$. Recently, inspired by the empirical success of information theoretic-based machine learning, several works have attempted to extend the PF method through the theory of variational inference [6]. In [7], the conditional mutual information form of PF due to $I(Y;Z)=I(X;Z)-I(X;Z|Y)$ is adopted, whose variational surrogate bound is estimated with empirical samples. In [8] a similar formulation is adopted, but they implement a generative deep architecture for interchangeable use of the empirical and the generated $Y$ samples for estimation instead. However, owing to the dependency of the private information $Y$ in estimating $I(X;Z|Y)$, these previous solvers violate the fundamental Markov relation in PF method. Consequently, samples of $X,Y$ are required for their solvers which means supervision is required in both model training and inference phases.
@@
Different from the previous works, we focus on the computation efficiency of PF in both known and unknown (empirical samples available) joint distribution settings based on the difference-of-convex (DC) structure of (1). In known joint distribution settings, inspired by the difference-of-convex algorithm (DCA), we propose a novel efficient PF solver with closed-form update equation. In our evaluation, the proposed solver not only covers solutions obtained from the state-of-the-art solvers but also attains non-trivial solutions that the compared greedy solvers are infeasible to reach. In unknown $P(X,Y)$ settings, our DC method complies with the fundamental Markov relation, therefore requires public information $X$ only in inference phase in contrast to previous approaches. We implement our solver with deep neural networks and evaluate the model on MNIST and Fashion-MNIST datasets. In all evaluated datasets, our approach outperforms the state-of-the-art. Specifically, under a comparable reconstruction quality of the public information $X$, an adversary’s clustering accuracy (leakage of the private information $Y$) with the released codes $Z$ of ours is significantly lower than the compared approaches.
@@
Upper-case letters denote random variables (RVs) and lower-case letters are their realizations. The calligraphic letter denotes the sample space, e.g., $x\in\mathcal{X}$ with cardinality $|\mathcal{X}|$. The vectors and matrices are denoted in bold-face symbols, i.e., $\boldsymbol{x},\boldsymbol{X}$ respectively. We denote the conditional probability mass function as a long vector:
@@ II A Difference-of-Convex Algorithm for Privacy Funnel
For simplicity, we focus on discrete random variables settings in this section. We start with a summary of DC programming and refer to the excellent review [9] for the development and recent advances of the DC method. For a DC program, the canonical form is given by:
@@
where $f,g$ are functions of $w$, and $\mathcal{W}$ denotes the feasible solution set of $w$. (4) relates to the PF method because (1) can be re-written into an unconstrained PF Lagrangian:
@@
where $\beta>0$ is a multiplier. When $P(X,Y)$ is known, the above reveals a DC structure. This is because $I(Z;X)$ is convex w.r.t. $P(Z|X)$ for known $P(X)$, and $-H(Z|Y)$ is convex w.r.t. $P(Z|X)$ as the Markov relation $P(Z|Y)=\sum_{x\in\mathcal{X}}P(Z|x)P(x|Y)$ along with known $P(Y)$ implies convexity [2]. The DC structure of (5) can be leveraged for efficient optimization. The idea is to use the first order approximation of the concave part of the objective function, resulting in a convex sub-objective. Then we iteratively minimize the convex surrogate sub-objective until the loss saturates. The monotonically decreasing loss values is assured due to the convexity of the sub-objective functions but the optimality of the converged solution is lost due to the non-convexity of the overall objective. There are well-developed efficient solvers for DC problems in literature, known as the Difference-of-Convex Algorithm (DCA). However, one of the key challenges in solving DC problems with DCA is that there exist infinitely many combinations of DC pairs [9]. Furthermore, a poorly chosen pair may result in slow convergence and numerical instability. We address it by proposing the following DC separation:
@@
where $\boldsymbol{B}_{y|x}$ denotes the matrix-form of $P(Y|X)$ with the $(i,j)$-entry represents $P(y_{i}|x_{j})$. We follow the common assumption in private source coding literature and assume that $\boldsymbol{B}_{y|x}$ is full-row rank [10, 11, 12, 13, 14]. Under this assumption, the pseudo-inverse of $\boldsymbol{B}_{y|x}$ exists, and we denote its Moore-Penrose pseudo inverse as $\boldsymbol{B}^{\dagger}_{y|x}:=(\boldsymbol{B}_{y|x}^{T}\boldsymbol{B}_{y|x})^{-1}\boldsymbol{B}_{y|x}^{T}$. Then (9) reduces to:
@@
where $\boldsymbol{c}^{k}_{z|x}:=\log{\boldsymbol{p}^{k}_{z}}+\beta\log{\frac{\boldsymbol{p}^{k}_{z|x}}{\boldsymbol{p}^{k}_{z}}}$, $\tilde{\boldsymbol{p}}_{z|y}$ denotes the un-normalized probability vector and $\propto$ represents proportional to. After normalization, followed by the Markov relation, we have $\boldsymbol{A}_{x|y}\boldsymbol{p}_{z|x}=\boldsymbol{p}_{z|y}$ with $\boldsymbol{A}_{x|y}$ denotes the matrix-form of $P(X|Y)$. Therefore, the insight from applying DCA to PF gives an update equation:
@@
where the $z\in\mathcal{Z}$ element is given by $\phi_{z}=e^{z}/\sum_{z^{\prime}\in\mathcal{Z}}e^{z^{\prime}}$. In practice, $\boldsymbol{\phi}$ is implemented as a softmax function. The above implies that the update of $\boldsymbol{p}_{z|x}$ from step $k$ to step $k+1$ requires solving the linear equation (11). However, since $\boldsymbol{B}_{y|x}$ is only full-row rank, (11) is an under-determined linear program over non-negative simplice (convex set). Heuristically, one would prefer a solution with the lowest complexity, which is often achieved by imposing $q$-norm (e.g., $q\in\{0,1,2\}$) constraints as regularization [15]. Based on this intuition, we therefore state the iterative update equation as the following optimization problem:
@@
where $\lVert\boldsymbol{v}\rVert_{q}:=(\sum_{v\in\mathcal{V}}|v|^{q})^{1/q}$. Similar to relaxing the “hard” constraints in classical regression problems, which allows for the application of significantly more efficient algorithms, we solve the following relaxed problem of (12) where a soft constraint is considered instead:
@@
where $\alpha>0$ is a relaxation coefficient. It is straightforward to see that the optimization problem is relaxed because of the tolerance of an approximation error of the linear equation. Remarkably, (12) is closely related to classical regression problems such as Ridge regression [16], LASSO [17], and sparse recovery [18]. But the key difference is that in (12) and (13), the feasible solution set $\Omega_{z|x}$ is the probability simplice. For $q=2$, since here (13) corresponds to a convex optimization problem, each step can be solved efficiently with off-the-shelf ridge regression solvers [19]. As for $q=1$ (convex surrogate for $q=0$), however, the LASSO solver [17] does not work well. This is because $\lVert\boldsymbol{p}_{z|x}\rVert_{1}=\sum_{x\in\mathcal{X}}|\boldsymbol{p}_{Z|x}|=|\mathcal{X}|$, since $\sum_{z\in\mathcal{Z}}p(z|x)=1,\forall x\in\mathcal{X}$ always. To address this, we provide an alternative approach for sparse recovery ($q=0$) with details referred to Appendix B:
@@
where $l_{x|y}:=\log{\boldsymbol{p}_{x|y}}$, $lse(x):=\log\sum_{x\in\mathcal{X}}\exp\{x\}$ is the log-sum-exponential function; $\mathcal{H}_{z|x}$ is the feasible set for $l_{z|x}\in[-M,-m]$ for some $M>m>0$ such that $\sum_{z\in\mathcal{Z}}\exp\{l_{z|x}\}=1,\forall x\in\mathcal{X}$; the subscript denotes the variable for summation, and $l_{\phi}^{k}:=\log{\boldsymbol{\phi}(\boldsymbol{B}_{y|x}^{\dagger}\boldsymbol{c}^{k}_{z|x})}$. Then after solving (14), we can project the obtained $l_{z|x}$ to a feasible solution $\boldsymbol{p}_{z|x}$ through a softmax function: $\boldsymbol{p}_{z|x}=\text{Softmax}(l_{z|x})$.
@@
For both implementations (13) ($q=2$) and (14), the sequence $\{\boldsymbol{p}_{z|x}^{k}\}_{k\in\mathbb{N}}$, obtained from (8), converges to a stationary point $p_{z|x}^{*}$ such that $\nabla f(\boldsymbol{p}^{*}_{z|x})=g(\boldsymbol{p}^{*}_{z|x})$ with $f,g$ defined in (6a).
@@ III Extension to Unknown Distributions
The main difficulty in applying PF is the full knowledge of the joint distribution $P(X,Y)$ [1]. In (13), this corresponds to defining the operators $A_{x|y}$ and $B_{y|x}^{\dagger}$. Without knowing $P(X,Y)$, these operators are intractible. To address this, we leverage the DC structure of the PF problem again, but now take expectation with respect to $P(Z,X)=P(Z|X)P(X)$ on the update equation which reveals that (see Appendix C):
@@
where the superscript $k$ denotes the iteration counter. The above update equation implies that the step $k+1$ solution $P^{k+1}(Z|X)$ is obtained from solving (15), given the previous step $k$ solution $P^{k}(Z|X)$.
@@
where the above surrogate bound is tight when $Q_{\phi}(Y|Z)=P(Y|Z)$. As for the other terms of (15), we parameterize $P_{\theta}(Z)=\mathcal{N}(\boldsymbol{0},\text{diag}(\boldsymbol{\sigma}_{\theta}^{2}))$ as standard Gaussian and $P_{\phi}(X|Z)=\mathcal{N}(\boldsymbol{\mu}_{\phi}(\boldsymbol{z}),\boldsymbol{I})$ as conditional Gaussian distributions. This transforms the intractible problem into a standard parameter estimation. Also, we parameterize the encoder $P_{\theta}(Z|X)$, corresponding to the expectation operator in (15), through the evident lower bound technique as adopted in the variational autoencoders (VAE) [20]. Here, we have $P_{\theta}(Z|X)\sim\mathcal{N}(\boldsymbol{\mu}_{\theta}(\boldsymbol{x}),\boldsymbol{\Sigma}_{\theta}(\boldsymbol{x}))$. Combining the above, we propose the following alternating solver:
$\phi^{*}:=\underset{\phi\in\Phi}{\arg\min}\,$ $-\beta\mathbb{E}_{x,z;\theta}\left[\log{P_{\phi}(X|Z)}\right]$ (17a)
$-\mathbb{E}_{y,z}\left[\log{Q_{\phi}(Y|Z)}\right],$
$\theta^{k+1}:=\underset{\theta\in\Theta}{\arg\min}\,$ $\frac{1}{2}\lVert I_{\phi^{*}}(Z;Y)+D_{KL}[P_{z;\theta}\parallel P^{(k)}_{z;\theta}]$ (18a)
@@
where $r_{Z}=\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ is a reference probability density function, regularizing the encoder; and $\alpha>0$ is a control threshold. Essentially, we alternate between fitting the empirical samples of $X,Y$ (17a) and solve the DCA update sub-problem with an extra regularization term (18a). The two steps are repeated until a pre-determined number of iterations is reached.
@@
  • While we adopt the variational inference similar to the previous works [7, 8]. We apply it on the expectation form of the proposed DC update equation (15), not on the PF Lagrangian nor on $I(X;Z|Y)$.

### v3

@@ I Introduction
where $\eta_{X}$ denotes a control threshold. In the above problem, the joint distribution $P(X,Y)$ is assumed to be known and the random variables satisfy the Markov chain $Y\rightarrow X\rightarrow Z$. Here, $Y$ is the private information; $X$ the public information and $Z$ is the released (compressed) codes; $I(Z;X)$ denotes the mutual information between two random variables $Z$ and $X$. In (1), the variables to optimize with are the stochastic mappings $P(Z|X)$ with $\Omega_{Z|X}$ denotes the associated feasible solution set. The aim of the problem is to characterize the fundamental trade-off between the privacy leakage and the utility of the released information. One of the challenges in solving (1) is the non-convexity of the problem. It is well-known that the mutual information $I(Z;X)$ with known marginal $P(X)$ is a convex function with respect to $P(Z|X)$ [2]; Additionally, due to the Markov relation, $I(Z;Y)$ is a convex function of $P(Z|X)$ as well. Together, (1) minimizes a convex objective under a non-convex set. Previous solvers for PF address the non-convexity by restricting the feasible solution set further for efficient computation. Both [1] and [3] consider discrete settings and focus on deterministic $P(z|x)$, i.e., a realization of $x\in\mathcal{X}$ belongs to a single cluster $z\in\mathcal{Z}$ only. They propose greedy algorithms to solve the reduced problem and differ in choosing the subset of $\mathcal{Z}$ to combine with. However, the imposed restriction limits the characterization of the privacy-utility trade-off. In our earlier work [4], we propose a non-convex alternating direction methods of multipliers (ADMM) solver for PF which is non-greedy, but it suffers from slow convergence [5].
@@
Another challenge of PF is knowing the joint distribution $P(X,Y)$. Overcoming this is important for practical application of the method as $P(X,Y)$ is prohibitively difficult to obtain in general. Nonetheless, it is relatively easy to obtain samples of the joint distribution, which become a training dataset as an indirect access to $P(X,Y)$. Recently, inspired by the empirical success of information theoretic-based machine learning, several works have attempted to extend the PF method through the theory of variational inference [6]. In [7], the conditional mutual information form of PF due to $I(Y;Z)=I(X;Z)-I(X;Z|Y)$ is adopted, whose variational surrogate bound is estimated with empirical samples. In [8] a similar formulation is adopted, but they restrict the parameter space and implement a generative deep architecture for interchangeable use of the empirical and the generated $Y$ samples for estimation. However, owing to the dependency of the private information $Y$ in estimating $I(X;Z|Y)$, these previous solvers violate the fundamental Markov relation in PF method. Consequently, samples of $X,Y$ are required for their solvers which means supervision is required in both model training and inference phases.
@@
Different from the previous works, we focus on more computationally efficient PF solvers for both known and unknown (empirical samples available) joint distribution settings based on the difference-of-convex (DC) structure of (1). In known joint distribution settings, inspired by the difference-of-convex algorithm (DCA), we propose a novel PF solver with closed-form update equation. In our evaluation, the proposed solver not only covers solutions obtained from the state-of-the-art solvers but also attains non-trivial solutions that the compared greedy solvers are infeasible to reach. In unknown $P(X,Y)$ settings, our DC method complies with the fundamental Markov relation, therefore requires public information $X$ only in inference phase in contrast to previous approaches. We implement our solver with deep neural networks and evaluate the model on MNIST and Fashion-MNIST datasets. In all evaluated datasets, our approach outperforms the state-of-the-art. Specifically, under a comparable reconstruction quality of the public information $X$, an adversary’s clustering accuracy (leakage of the private information $Y$) with the released codes $Z$ of ours is significantly lower than the compared approaches.
@@
Upper-case letters denote random variables (RVs) and lower-case letters are their realizations. The calligraphic letter denotes the sample space, e.g., $x\in\mathcal{X}$ with cardinality $|\mathcal{X}|$. The vectors and matrices are denoted in bold-face symbols, e.g., $\boldsymbol{x},\boldsymbol{X}$ respectively. We denote the conditional probability mass function as a long vector:
@@ II A Difference-of-Convex Algorithm for Privacy Funnel
For simplicity, we focus on discrete random variables settings in this section. We start with a summary of DC programming and refer to the excellent review [9] for the development and recent advances. For a DC program, the canonical form is given by:
@@
where $f,g$ are functions of $w$, and $\mathcal{W}$ denotes the feasible solution set of $w$. (3) relates to the PF method through rewriting (1) as an unconstrained PF Lagrangian:
@@
where $\beta>0$ is a multiplier. The above is a DC program when $P(X,Y)$ is known for a fixed $\beta$. This is because $I(Z;X)$ is convex w.r.t. $P(Z|X)$ for known $P(X)$, and $-H(Z|Y)$ is convex w.r.t. $P(Z|X)$ as the Markov relation $P(Z|Y)=\sum_{x\in\mathcal{X}}P(Z|x)P(x|Y)$ for known $P(Y)$ [2]. The DC structure of (4) can be leveraged for efficient optimization. By substituting the concave part with its first order approximation, we obtain a convex sub-objective. Then we iteratively minimize the convex surrogate sub-objective until the loss saturates. The monotonically decreasing loss values is assured due to the convexity of the sub-objective functions. However, the optimality of the converged solution is lost due to the non-convexity of the overall problem. There are well-developed efficient solvers for DC problems in literature, known as the Difference-of-Convex Algorithm (DCA). However, one of the key challenges in solving DC problems with DCA is that there exist infinitely many combinations of DC pairs [9]. Furthermore, a poorly chosen pair may result in slow convergence and numerical instability. We address the issues by proposing the following DC separation:
@@
where $\boldsymbol{B}_{y|x}:=\boldsymbol{I}_{|\mathcal{Z}|}\otimes\boldsymbol{Q}_{y|x}$, with $\boldsymbol{I}_{d}$ denotes the $d$-dimensional identity matrix, $\otimes$ the Kronecker product, and $\boldsymbol{Q}_{y|x}$ the matrix-form of $P(Y|X)$ whose $(i,j)$-entry is $P(y_{i}|x_{j})$. We follow the assumption in private source coding literature that $\boldsymbol{B}_{y|x}$ is full-row rank [10, 11, 12, 13, 14]. Then we denote the Moore-Penrose pseudo-inverse of $\boldsymbol{B}_{y|x}$ as $\boldsymbol{B}^{\dagger}_{y|x}:=(\boldsymbol{B}_{y|x}^{T}\boldsymbol{B}_{y|x})^{-1}\boldsymbol{B}_{y|x}^{T}$. Note that due to the Markov relation, we have $\boldsymbol{A}_{x|y}\boldsymbol{p}_{z|x}=\boldsymbol{p}_{z|y}$ with $\boldsymbol{A}_{x|y}:=\boldsymbol{I}_{\mathcal{Z}}\otimes\boldsymbol{Q}_{x|y}$ defined similarly as $\boldsymbol{B}_{y|x}$. Summarizing the above, (8) reduces to:
@@
where $\boldsymbol{c}^{k}_{z|x}:=\log{\boldsymbol{p}^{k}_{z}}+\beta\log{\frac{\boldsymbol{p}^{k}_{z|x}}{\boldsymbol{p}^{k}_{z}}}$, $\tilde{\boldsymbol{p}}_{z|y}$ denotes the un-normalized probability vector and $\propto$ represents proportional to. After normalization, followed by the Markov relation, the insight from applying DCA to PF gives an update equation:
@@
where the $z\in\mathcal{Z}$ element is given by $\phi_{z}=e^{z}/\sum_{z^{\prime}\in\mathcal{Z}}e^{z^{\prime}}$. In practice, $\boldsymbol{\phi}$ is implemented as a softmax function. The above implies that the update of $\boldsymbol{p}_{z|x}$ from step $k$ to step $k+1$ requires solving the linear equation (10). However, since $\boldsymbol{A}_{x|y}^{\dagger}$ might not exist ($\boldsymbol{B}_{y|x}$ is full-row rank), (10) is an under-determined linear program over non-negative simplice. Heuristically, one would prefer a solution with the lowest complexity, which is often achieved by imposing $q$-norm (e.g., $q\in\{0,1,2\}$) constraints as regularization [15]. Based on this intuition, we therefore solve the following sub-problem:
@@
where $\lVert\boldsymbol{v}\rVert_{q}:=(\sum_{v\in\mathcal{V}}|v|^{q})^{1/q}$. Similar to relaxing the “hard” constraints in classical regression problems for applying more efficient algorithms, we solve the following relaxed problem of (11) where a soft constraint is considered instead:
@@
where $\alpha>0$ is a relaxation coefficient. The relaxed problem tolerates an approximation error of the linear equation instead of a strict equality constraint. Remarkably, (11) relates to the classical regression problems such as Ridge regression [16], LASSO [17], and sparse recovery [18]. But in (11) and (12), the feasible solution set $\Omega_{z|x}$ is the probability simplice. For $q=2$, since (12) corresponds to a convex optimization problem, each step can be solved efficiently with off-the-shelf ridge regression solvers [19]. As for $q=1$ (convex surrogate for $q=0$), however, the LASSO solver [17] does not work well. This is because $\lVert\boldsymbol{p}_{z|x}\rVert_{1}=\sum_{x\in\mathcal{X}}|\boldsymbol{p}_{Z|x}|=|\mathcal{X}|$, since $\sum_{z\in\mathcal{Z}}P(z|x)=1,\forall x\in\mathcal{X}$ always. To address this, we provide the following alternative for sparse recovery ($q=0$) with details referred to Appendix B:
@@
where $l_{x|y}:=\log{\boldsymbol{p}_{x|y}}$, $lse(x):=\log\sum_{x\in\mathcal{X}}\exp\{x\}$ is the log-sum-exponential function; $\mathcal{H}_{z|x}$ is the feasible set for $l_{z|x}\in[-M,-m]$ for some $M>m>0$ such that $\sum_{z\in\mathcal{Z}}\exp\{l_{z|x}\}=1,\forall x\in\mathcal{X}$; the subscript of lse denotes the variable for summation, and $l_{\phi}^{k}:=\log{\boldsymbol{\phi}(\boldsymbol{B}_{y|x}^{\dagger}\boldsymbol{c}^{k}_{z|x})}$. Then after solving (13), we can project the obtained $l_{z|x}$ to a feasible solution $\boldsymbol{p}_{z|x}$ through a softmax function: $\boldsymbol{p}_{z|x}=\text{Softmax}(l_{z|x})$.
@@
For both implementations (12) ($q=2$) and (13), the sequence $\{\boldsymbol{A}_{x}\boldsymbol{p}_{z|x}^{k}\}_{k\in\mathbb{N}}$, obtained from (7), converges to a stationary point $\boldsymbol{A}_{x}\boldsymbol{p}_{z|x}^{*}$ such that $\nabla f(\boldsymbol{p}^{*}_{z|x})=\nabla g(\boldsymbol{p}^{*}_{z|x})$ with $f,g$ defined in (5a), and $\boldsymbol{A}_{x}:=\boldsymbol{I}_{\mathcal{Z}}\otimes\boldsymbol{Q}_{x}$.
@@ III Extension to Unknown Distributions
The main difficulty in applying PF is the full knowledge of the joint distribution $P(X,Y)$ [1]. In (12), this corresponds to defining the operators $\boldsymbol{A}_{x|y}$ and $\boldsymbol{B}_{y|x}^{\dagger}$. Without knowing $P(X,Y)$, these operators are intractible. To address this, we leverage the DC structure of the PF problem again, but now take expectation with respect to $P(Z,X)=P(Z|X)P(X)$ on the update equation (see Appendix C):
@@
where the superscript $k$ denotes the iteration counter. The above update equation implies that the step $k+1$ solution $P^{k+1}(Z|X)$ is obtained from solving (14), given the step $k$ solution $P^{k}(Z|X)$.
@@
where the above surrogate bound is tight when $Q_{\phi}(Y|Z)=P(Y|Z)$. As for the other terms in (14), we parameterize $P_{\theta}(Z)=\mathcal{N}(\boldsymbol{0},\text{diag}(\boldsymbol{\sigma}_{\theta}^{2}))$ as standard Gaussian and $P_{\phi}(X|Z)=\mathcal{N}(\boldsymbol{\mu}_{\phi}(\boldsymbol{z}),\boldsymbol{I})$ as conditional Gaussian distributions. This transforms the intractible problem into standard parameter estimation. Also, we parameterize the encoder $P_{\theta}(Z|X)$, corresponding to the expectation operator in (14), through the evident lower bound technique as adopted in the variational autoencoders (VAE) [20]. Here, we have $P_{\theta}(Z|X)\sim\mathcal{N}(\boldsymbol{\mu}_{\theta}(\boldsymbol{x}),\boldsymbol{\Sigma}_{\theta}(\boldsymbol{x}))$. Combining the above, we propose the following alternating solver:
$\phi^{*}:=\underset{\phi\in\Phi}{\arg\min}\,$ $-\beta\mathbb{E}_{x,z;\theta^{k}}\left[\log{P_{\phi}(X|Z)}\right]$ (16a)
$-\mathbb{E}_{y,z;\theta^{k}}\left[\log{Q_{\phi}(Y|Z)}\right],$
$\theta^{k+1}:=\underset{\theta\in\Theta}{\arg\min}\,$ $\frac{1}{2}\lVert I_{\phi^{*};\theta}(Z;Y)+D_{KL}[P_{z;\theta}\parallel P_{z;\theta^{k}}]$ (17a)
@@
where $r_{Z}=\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ is a reference probability density function, regularizing the encoder; and $\alpha>0$ is a control threshold. Essentially, we alternate between fitting the empirical samples of $X,Y$ (16a) and solving the DCA update sub-problem with an extra regularization term (17a). The two steps are repeated until a pre-determined number of iterations is reached.
@@
  • While variational inference is adopted of the proposed solver, but we apply it on the expectation form of the proposed DC update equation (14), not on the PF Lagrangian nor on $I(X;Z|Y)$ [7, 8].